A certain forest covers an area of 2000 km^2 . Suppose that each year this area decreases by 4.25%. What will the area be after 6 years?

it asks for the answer in ^2km.
What I think i should do is order of operations
2000^2 *4.25% *6
2000^2×425×6
4000000×425×6
1700000000×6
10200000000
I feel like this is wrong because it ask for the answer in ^2 km

Answer by jim_thompson5910(35256)

A = P(1-r)^t

A = 2000(1-0.0425)^t

A = 2000(0.9575)^t

A = 2000(0.9575)^6

A = 2000(0.7706065)

A = 1541.213

So the answer is roughly 1541.213 square km